Lance hiked up a mountain at a rate of 2.5 mph and returned along the same trail at 3mph. His total hiking time was 11 hours. How long was the hiking trail?
Answer by rwm(914)


time up =p
time down=n
distance up =d
distance down=d
p+n=11
r*t=d
2.5*p=d
3*n=d
d=15 hiking trail
n=5
p=6